ms.noemi jogged at a total distance mof 6600 yards in threee nights.If each night she incr5eased her distance to 440 yards,how far did she jog on the first night.
Found 2 solutions by stanbon, BurningFlame_marc:
Answer by stanbon(75887)

ms.noemi jogged at a total distance mof 6600 yards in threee nights.If each night she incr5eased her distance to 440 yards,how far did she jog on the first night.
----------
1st night distance: x
2nd night distance:x+440
3rd night distance: x+440+440 = x+880
--------------------
Equation:
x + x+440 + x+880 = 6600
3x + 1320 = 6600
3x = 5280
x = 1760 yrds.(1st night distance)
==================================
Cheers,
Stan H.

Answer by BurningFlame_marc(77)

let x=distance she jogged in the 1st night; x+440=distance jogged in the 2nd night; x+880=distance jogged in the 3rd nigth.
x+(x+440)+(x+880)=6600
3x+1320=6600
3x=6600-1320
3x=5280
x=1760
she jogged 1760 yards on the first night.